OpenStudy (anonymous):

hhehhhhhhlllllppppp.......... Add the following polynomials, then place the answer in the proper location on the grid. 2x^3-4x^2+6x-3 x^3+6x^2-8x+12 -3x^3+2x^2-4x-7 ____________________

OpenStudy (anonymous):

are you trying to find the degree

OpenStudy (anonymous):

no, the polynomials...

OpenStudy (anonymous):

i don't think you can factor these

OpenStudy (anonymous):

well it is one of my question in my algebra.....i am suppose to add them or something..

OpenStudy (anonymous):

it say you are suppose to but the answer in the grid. thats why I asked you about the degree.

OpenStudy (anonymous):

put :/

OpenStudy (anonymous):

The degree is the sum of the exponents of each variable in the expression

OpenStudy (anonymous):

i do not follow...:( sorry???

OpenStudy (anonymous):

you have to place your answer on the grid? Yes?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

ok so now we are on the same page :)

OpenStudy (anonymous):

2x^3-4x^2+6x-3 The degree is the sum of the exponents of each variable in the expression. In this case, the degree of 2x^3 = 3 -4x^2=2 -3 = 0 The degree of 2x^3-4x^2+6x-3 is the largest degree of its terms so it would be 3

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

that i not all is it????

OpenStudy (anonymous):

they are all 3

OpenStudy (anonymous):

wait let me check something

OpenStudy (anonymous):

they gave me this hint ____^2_____________

OpenStudy (anonymous):

well individually the all are three but if you add up all 3 of the equations it comes out to 2 sorry I might not be sure what you are asking

OpenStudy (anonymous):

if the question reads place the answer thaen I would assume that there is only one answer so if you add up all three equations the answer is 2 If you do them individually they all have the same answer 3

OpenStudy (anonymous):

hope that helps